OpenStudy (anonymous):

what is the exact value of sin 45 degrees as found on the unit circle

OpenStudy (mikeymaximum):

@Nerd94

OpenStudy (anonymous):

25

OpenStudy (anonymous):

that doesn't make any sense it needs to be in terms of pi

OpenStudy (ganpat):

unit circle means radius = 1 units..  sin 45 = x/r (opp/hyp) r2 = x2 + y2 given, r = 1 and x = y 1 = x2 + y2 1 = 2x2 x = 1/sq r t (2) sin 45 = 1/sq rt (2)

OpenStudy (ganpat):

i am not sure, this is what i can recall :D :p

OpenStudy (anonymous):

i wrote this wrong it was supposed to be cos instead of sin… sooooo sorry!! @Ganpat

OpenStudy (ganpat):

ha ha haa. can u calculate your self using above formulas ?? not much will change..

OpenStudy (anonymous):

i get mixed up between what a and o are

OpenStudy (ganpat):

huh?? there is no a and o !!  anyways, we have formed a right angle triangle inside the circle.. SO angle b has to be 45 degree.. ( 180= 90 + 45 + 45) also it is a unit circle.. so r = 1.. Now find cos O